# DEPF: A UAV Multispectral Object Detector with Dual-Domain Enhancement and Priority-Guided Mamba Fusion

### v1

@@ Introduction
To evaluate our priority scanning, the visualization of fusion feature maps are shown in Figure 6. It could be observed that the degree of focus is located on background easily under Z-order fusion mechanism. By contrast, our priority scanning fusion shifts the focus of attention to the instances, suppressing the interference from background information.
Other Experiments
We add other experiment results in the appendix file of the supplementary materials.
Conclusion
@@
In this article, DEPF is proposed with DDE module and PGMF module. DDE contains CSWM and FDR block to enhance the global brightness and recover texture-details of low-light RGB images. PGMF introduces Priority Scanning to enhance the local target modeling and achieve feature alignment. To our best knowledge, PGMF is the first approach to improve Mamba mechanism for multi-modality fusion. The experiment results shows that DEPF performs well on object detection and achieves lightweight computation.

### v2

@@ IV-D3 Visualization of fusion feature maps
To evaluate our priority scanning, the visualization of fusion feature maps are shown in Fig 8. It could be observed that the degree of focus is located on background easily under Z-order fusion mechanism. By contrast, our priority scanning fusion shifts the focus of attention to the instances, suppressing the interference from background information.
IV-D4 Ablation Study on Fusion Ways
Fig VIII shows four different ways of fusion priority sequences. Additionally, in the Method Section, we have proved that the fusion sequence satisfies the format of Fig. 9 (d). Therefore, the ablation study about the fusion ways is conducted, whose results could be seen in Table VIII. The results of way (b) and way (c) prove that the locations of high priority tokens have an impact on accuracy, which shows that the early tokens are more important than late tokens because the former make better parameter renewing of $A$ matrix if related information is provided. For way (a), when scanning arrives at the end of sequences, the contributions of high priority tokens are decayed. The parameter renewing of $A$ matrix is affected by early tokens which contain background information. These two factors lead to unsteady accuracy, like mAP@0.5:0.95 of 48.9%. Following way (d) from our proof, DEPFusion obtains mAP@0.5 with 78.9% and mAP@0.5:0.95 of 51.2%.
Notably, the theory of PGMF module is proved from the perspective of contribution’s limit. Therefore, for an infinitely long sequence, the advantage of priority-guided serialization may be obvious. However, the sequences in real scenes are finite length almost. Therefore, our future work would explore the theory’s scope of application further.
Figure 9: Different ways of fusion priority sequences. (a) means $\overline{x_{vp}}+x_{ip}$; (b) means $x_{vp}+x_{ip}$; (c) means $\overline{x_{vp}}+\overline{x_{ip}}$ and (d) means $x_{vp}+\overline{x_{ip}}$.
TABLE VIII: Ablation study result on different fusion ways.
Fusing Way mAP@0.5 mAP@0.5:0.95
$\overline{x_{vp}}+x_{ip}$ (a) 76.6 48.9
$x_{vp}+x_{ip}$ (b) 76.4 50.6
$\overline{x_{vp}}+\overline{x_{ip}}$ (c) 75.8 49.7
$x_{vp}+\overline{x_{ip}}$ (d) 78.9 51.2
V Conclusion
@@
In this article, DEPFusion is proposed with DDE module and PGMF module. DDE contains CSWM and FDR block to enhance the global brightness and recover texture-details of low-light RGB images. PGMF introduces Priority Scanning to enhance the local target modeling and achieve feature alignment. To our best knowledge, PGMF is the first approach to improve Mamba mechanism for multi-modality fusion. The experiment results shows that DEPFusion performs well on object detection and achieves lightweight computation.